Reports 1-1 of 1 Clear search Modify search
On-call intervention (General)
Oncall-system - 23:47 Friday 14 May 2021 (51796) Print this report
On-call intervention

The following report has been submitted to the On-call interface.

On-call events -> Injection System

Title: Fast unlock crysis

Author(s): spinicelli, chiummo

Called at: 18:30, 14-05-2021, by: Menzione
Remote intervention: Started: 18:30, 14-05-2021; Ended:
On-site intervention: Started: ; Ended:
Status: Unresolved
Operator when issue resolved: Menzione


In the late afternoon IMC and RFC started to suddenly and repetitively unlock, creating problems to the planned activities of the ISC team.
After a first analysis, it is clear that this problem is  different from the one reported during the night of between 12th and 13th May ( 51771 ). Indeed, that phenomenon was a slow decrease of the power due to mode competition caused by a mode-hope instability, that has been solved today.

Instead, analyzing the raw data of some of this evening unlocks (plot 1, 2, 3) we can observe typical behaviour of EOM corrections saturation that triggers  the unlock of IMC and RFC. After a quite long period of instability between ~17:30 and 19:30 UTC, those unlocks started to calm down a little.

Up to now, it's unclear what could trigger these crysis. We leave the system with all loops closed.

* Note that any files attached to this report are available in the On-call interface.

Comments to this report:
spinicelli - 23:53 Friday 14 May 2021 (51797) Print this report

Plot 1-3 

Images attached to this comment
derossi, spinicelli, chiummo - 13:12 Monday 17 May 2021 (51810) Print this report

This morning we were called by Berni due to frequent IMC unlocks.

As an attempt to mitigate the problem we performed the following actions:

- we moved the electronics of the ML away from the same breadbord of the optical setup (1st attached picture).

- we tuned the input ML polarization by turning the waveplate before the collimator from 182° to 185° (2nd plot) while checking the stability when moving the fiber. By turning the waveplate we noticed that the output power dropped a lot on both polarziation. This behaviour could be due to the fiber connector between the two PM fibers.

- we installed a beam dump on the SL input in order to get rid from possible backreflection when SL unlocks occur.

- we finally lowered the PSTAB gain in order to have the usual UGF of 75 kHz (3rd picture). We had in fact more than 100 kHz.

We relocked everything and we will keep monitoring hte system.

Images attached to this comment
Search Help